# SAM2Long: Enhancing SAM 2 for Long Video Segmentation with a Training-Free Memory Tree

### v1

@@ Abstract
The Segment Anything Model 2 (SAM 2) has emerged as a powerful foundation model for object segmentation in both images and videos, paving the way for various downstream video applications. The crucial design of SAM 2 for video segmentation is its memory module, which prompts object-aware memories from previous frames for current frame prediction. However, its greedy-selection memory design suffers from the “error accumulation” problem, where an errored or missed mask will cascade and influence the segmentation of the subsequent frames, which limits the performance of SAM 2 toward complex long-term videos. To this end, we introduce SAM2Long, an improved training-free video object segmentation strategy, which considers the segmentation uncertainty within each frame and chooses the video-level optimal results from multiple segmentation pathways in a constrained tree search manner. In practice, we maintain a fixed number of segmentation pathways throughout the video. For each frame, multiple masks are proposed based on the existing pathways, creating various candidate branches. We then select the same fixed number of branches with higher cumulative scores as the new pathways for the next frame. After processing the final frame, the pathway with the highest cumulative score is chosen as the final segmentation result. Benefiting from its heuristic search design, SAM2Long is robust toward occlusions and object reappearances, and can effectively segment and track objects for complex long-term videos. Without introducing any additional parameters or further training, SAM2Long significantly and consistently outperforms SAM 2 on five VOS benchmarks. Notably, SAM2Long achieves an average improvement of 3.0 points across all 24 head-to-head comparisons, with gains of up to 5.3 points in $\mathcal{J}\&\mathcal{F}$ on long-term video object segmentation benchmarks such as SA-V and LVOS. The code is released at https://github.com/Mark12Ding/SAM2Long.
@@ 1 Introduction
Figure 1: Comparison of occlusion handling and long-term compatibility between SAM 2 and SAM2Long. (a) When an occlusion occurs, SAM 2 may lose track or follow the wrong object, leading to accumulated errors. In contrast, SAM2Long utilizes memory tree search to recover when the object reappears. (b) The per-frame $\mathcal{J}\&\mathcal{F}$ scores of the predicted masks are plotted at specific timestamps on the LVOS and SA-V datasets. SAM2Long demonstrates greater resilience to elapsed time compared to SAM 2, maintaining superior performance over longer periods.
@@
To this end, we redesign the memory module of SAM 2 to enhance its long-term compatibility and robustness against occlusions and error propagation. Our improvement is completely free of additional training and does not introduce any external parameters, but simply unleashes the potential of SAM 2 itself. Our approach is motivated by the observation that the SAM 2 mask decoder generates multiple diverse masks, accompanied by predicted IoU scores and an occlusion score when handling challenging and ambiguous cases. However, SAM 2 only selects a single mask as memory, sometimes disregarding the correct one. To address this, we aim to equip SAM 2 with multiple memory pathways, allowing various masks to be stored as memory at each time step, thereby improving predictions for subsequent frames.
@@ 2.1 Video Object Segmentation
Perceiving the environment in terms of objects is a fundamental cognitive ability of humans. In computer vision, Video Object Segmentation (VOS) tasks aim to replicate this capability by requiring models to segment and track specified objects within video sequences. A substantial amount of research has been conducted on video object segmentation in recent decades (Fan et al., 2019; Oh et al., 2019; Hu et al., 2018a; Oh et al., 2018; Perazzi et al., 2017; Wang et al., 2019; Hu et al., 2018b; Li & Loy, 2018; Bao et al., 2018; Zhang et al., 2019; Li et al., 2020; Johnander et al., 2019; Zhang et al., 2023; Ventura et al., 2019; Li et al., 2022; Wu et al., 2023; Wang et al., 2023; Qian et al., 2023; 2025; Ding et al., 2022; 2023b).
@@ 2.3 Segment Anything Model
Segment Anything Model (SAM) (Kirillov et al., 2023) is recognized as a milestone vision foundation model that can segment any object in an image using interactive prompts. Its impressive zero-shot transfer performance has shown great versatility in various vision tasks, including segmentation applications (Li et al., 2023; Ma et al., 2024; Xu et al., 2024), image editing (Gao et al., 2023) and object reconstruction (Lin et al., 2024).
@@ 4.2 Main Results
SAM2Long consistently improves SAM 2 over all model sizes and datasets. Table 1 presents an overall comparison between SAM 2 and SAM2Long across various model sizes on the SA-V validation and test sets, as well as the LVOS v2 validation set. In total, the table includes 8 model variations, covering SAM 2 and the latest SAM 2.1 across four model sizes. The average performance across 24 experiments yields a $\mathcal{J}\&\mathcal{F}$ score of 3.0. These results confirm that SAM2Long consistently outperforms the SAM 2 baseline by a large margin. For instance, for SAM2Long-Large, achieves an improvement of 4.5 and 5.3 over SAM 2 on SA-V val and test sets. This pattern is also evident in the LVOS validation set, where SAM2Long delivers notable performance gains over SAM 2 for each model size. These results showcase the effectiveness of the training-free memory tree in long-time video segmentation scenarios.
@@
SAM2Long demonstrates versatility when handling videos with various challenges. In addition to the SA-V and LVOS datasets, we evaluate our proposed SAM2Long on other VOS benchmarks in Table 4. On the MOSE dataset (Ding et al., 2023a), which involves complex real-world scenes, SAM2.1Long achieves a $\mathcal{J}\&\mathcal{F}$ score of 75.2, surpassing SAM 2.1’s score of 74.5. Given that the stronger SAM 2.1-L shows no improvement over the SAM 2-L model on the MOSE benchmark, our performance gain with SAM2.1Long is particularly notable. Similarly, on the VOST dataset (Tokmakov et al., 2023), which focuses on objects undergoing extreme transformations, SAM2.1Long shows improvement with a $\mathcal{J}\&\mathcal{F}$ score of 54.0, a nearly 1-point gain over SAM2.1. On the PUMaVOS dataset (Bekuzarov et al., 2023), which challenges models with difficult visual cues, SAM2.1Long outperforms SAM 2.1 with a score of 82.4 compared to 81.1, demonstrating its enhanced ability to handle subtle and ambiguous segmentation tasks. These results underscore that we have preserved the fundamental segmentation capabilities of SAM 2 while enhancing its long-term abilities, demonstrating the robustness and versatility of SAM2Long across a range of VOS benchmarks.
@@
Table 4: The performance comparisons between SAM 2 and SAM2Long on other VOS benchmarks. All experiments use SAM2.1-Large model.
Dataset SAM 2.1 SAM2.1Long
[…]
MOSE (Ding et al., 2023a) 74.5 70.6 78.4 75.2 71.1 79.3
VOST (Tokmakov et al., 2023) 53.1 47.8 58.3 54.0 48.4 59.6
PUMaVOS (Bekuzarov et al., 2023) 81.1 78.5 83.7 82.4 79.6 85.1
@@ 4.3 Ablation Study
In terms of speed, since we maintain a fixed number of memory pathways at every time step, the processing speed remains efficient. Using three memory pathways slows down the model by only $18\%$, while yielding a 4.5-point increase in performance.
@@
1 76.3 73.0 79.5 1$\times$
2 80.1 76.7 83.5 0.93$\times$
3 80.8 77.1 84.5 0.82$\times$
4 80.7 77.0 84.5 0.75$\times$
@@ 5 Conclusion
In this paper, we introduce SAM2Long, a training-free enhancement to SAM 2 that alleviates its limitations in long-term video object segmentation. By employing a constrained tree memory structure with object-aware memory modulation, SAM2Long effectively mitigates error accumulation and improves robustness against occlusions, resulting in a more reliable segmentation process over extended periods. Extensive evaluations on five VOS benchmarks demonstrate that SAM2Long consistently outperforms SAM 2, especially in complex video scenarios. Notably, SAM2Long achieves up to a 5-point improvement in $\mathcal{J}\&\mathcal{F}$ scores on challenging long-term video benchmarks such SA-V and LVOS without requiring additional training or external parameters.

### v2

# Shuangrui Ding1  Rui Qian1  Xiaoyi Dong1,2  Pan Zhang2 Yuhang Zang2  Yuhang Cao2  Yuwei Guo1   Dahua Lin1   Jiaqi Wang2

@@ Abstract
The Segment Anything Model 2 (SAM 2) has emerged as a powerful foundation model for object segmentation in both images and videos, paving the way for various downstream video applications. The crucial design of SAM 2 for video segmentation is its memory module, which prompts object-aware memories from previous frames for current frame prediction. However, its greedy-selection memory design suffers from the “error accumulation” problem, where an errored or missed mask will cascade and influence the segmentation of the subsequent frames, which limits the performance of SAM 2 toward complex long-term videos. To this end, we introduce SAM2Long, an improved training-free video object segmentation strategy, which considers the segmentation uncertainty within each frame and chooses the video-level optimal results from multiple segmentation pathways in a constrained tree search manner. In practice, we maintain a fixed number of segmentation pathways throughout the video. For each frame, multiple masks are proposed based on the existing pathways, creating various candidate branches. We then select the same fixed number of branches with higher cumulative scores as the new pathways for the next frame. After processing the final frame, the pathway with the highest cumulative score is chosen as the final segmentation result. Benefiting from its heuristic search design, SAM2Long is robust toward occlusions and object reappearances, and can effectively segment and track objects for complex long-term videos. Without introducing any additional parameters or further training, SAM2Long significantly and consistently outperforms SAM 2 on six VOS benchmarks. Notably, SAM2Long achieves an average improvement of 3.0 points across all 24 head-to-head comparisons, with gains of up to 5.3 points in $\mathcal{J}\&\mathcal{F}$ on long-term video object segmentation benchmarks such as SA-V and LVOS. The code is released at https://github.com/Mark12Ding/SAM2Long.
@@ 1 Introduction
Figure 1: Comparison of occlusion handling and long-term capability between SAM 2 and SAM2Long. (a) When an occlusion occurs, SAM 2 may lose track or follow the wrong object, leading to accumulated errors. In contrast, SAM2Long utilizes memory tree search to recover when the object reappears. (b) The per-frame $\mathcal{J}\&\mathcal{F}$ scores of the predicted masks are plotted at specific timestamps on the LVOS and SA-V datasets. SAM2Long demonstrates greater resilience to elapsed time compared to SAM 2, maintaining superior performance over longer periods.
@@
To this end, we redesign the memory module of SAM 2 to enhance its long-term capability and robustness against occlusions and error propagation. Our improvement is completely free of additional training and does not introduce any external parameters, but simply unleashes the potential of SAM 2 itself. Our approach is motivated by the observation that the SAM 2 mask decoder generates multiple diverse masks, accompanied by predicted IoU scores and an occlusion score when handling challenging and ambiguous cases. However, SAM 2 only selects a single mask as memory, sometimes disregarding the correct one. To address this, we aim to equip SAM 2 with multiple memory pathways, allowing various masks to be stored as memory at each time step, thereby improving predictions for subsequent frames.
@@ 2.1 Video Object Segmentation
Perceiving the environment in terms of objects is a fundamental cognitive ability of humans. In computer vision, Video Object Segmentation (VOS) tasks aim to replicate this capability by requiring models to segment and track specified objects within video sequences. A substantial amount of research has been conducted on video object segmentation in recent decades (Fan et al., 2019a; Oh et al., 2019; Hu et al., 2018a; Oh et al., 2018; Perazzi et al., 2017; Wang et al., 2019; Hu et al., 2018b; Li & Loy, 2018; Bao et al., 2018; Zhang et al., 2019; Li et al., 2020; Johnander et al., 2019; Zhang et al., 2023; Ventura et al., 2019; Li et al., 2022; Wu et al., 2023b; Wang et al., 2023; Qian et al., 2023; 2025; Ding et al., 2022; 2023b).
@@ 2.3 Segment Anything Model
Segment Anything Model (SAM) (Kirillov et al., 2023) is recognized as a milestone vision foundation model that can segment any object in an image using interactive prompts. Its impressive zero-shot transfer performance has shown great versatility in various vision tasks, including segmentation applications (Li et al., 2023; Ma et al., 2024; Xu et al., 2024), image editing (Gao et al., 2023a) and object reconstruction (Lin et al., 2024).
@@ 4.2 Main Results
SAM2Long consistently improves SAM 2 over all model sizes and datasets. Table 1 presents an overall comparison between SAM 2 and SAM2Long across various model sizes on the SA-V validation and test sets, as well as the LVOS v2 validation set. In total, the table includes 8 model variations, covering SAM 2 and the latest SAM 2.1 across four model sizes. The average performance across 24 experiments shows an improvement of 3 points in the $\mathcal{J}\&\mathcal{F}$ score. These results confirm that SAM2Long consistently outperforms the SAM 2 baseline by a large margin. For instance, for SAM2Long-Large, achieves an improvement of 4.5 and 5.3 over SAM 2 on SA-V val and test sets. This pattern is also evident in the LVOS validation set, where SAM2Long delivers notable performance gains over SAM 2 for each model size. These results showcase the effectiveness of the training-free memory tree in long-time video segmentation scenarios.
@@
SAM2Long demonstrates versatility when handling videos with various challenges. In addition to the SA-V and LVOS datasets, we evaluate our proposed SAM2Long on other VOS benchmarks in Table 4. On the PUMaVOS (Bekuzarov et al., 2023), which challenges models with difficult visual cues, SAM2.1Long outperforms SAM 2.1 with a score of 82.4 compared to 81.1, demonstrating its enhanced ability to handle subtle and ambiguous segmentation tasks. Similarly, on the VOST (Tokmakov et al., 2023), which focuses on objects undergoing extreme transformations, SAM2.1Long shows improvement with a $\mathcal{J}\&\mathcal{F}$ score of 54.0, a nearly 1-point gain over SAM2.1. On the MOSE (Ding et al., 2023a), which involves complex real-world scenes, SAM2.1Long achieves a $\mathcal{J}\&\mathcal{F}$ score of 75.2, surpassing SAM 2.1’s score of 74.5. Given that the stronger SAM 2.1-L shows no improvement over the SAM 2-L model on the MOSE benchmark, our performance gain with SAM2.1Long is particularly notable. However, YouTube-VOS Xu et al. (2018), with an average duration of only 4.5 seconds, offers limited opportunities to showcase long-term tracking capabilities. Here, both SAM2.1 and SAM2.1Long achieve the same score of 88.7. These results underscore that we have preserved the fundamental segmentation capabilities of SAM 2 while enhancing its long-term abilities, demonstrating the robustness and versatility of SAM2Long across a range of VOS benchmarks.
@@
Dataset Duration SAM 2.1 SAM2.1Long
seconds $\mathcal{J}\&\mathcal{F}$ $\mathcal{J}\&\mathcal{F}$
PUMaVOS (Bekuzarov et al., 2023) 28.7 81.1 82.4
VOST (Tokmakov et al., 2023) 21 53.1 54.0
MOSE (Ding et al., 2023a) 12.4 74.5 75.2
YTVOS (Xu et al., 2018) 4.5 88.7 88.7
Table 4: The performance comparisons between SAM 2 and SAM2Long on other VOS benchmarks. We list the average duration for each benchmark. All experiments use SAM2.1-L model.
@@ 4.3 Ablation Study
In terms of speed, since we maintain a fixed number of memory pathways at every time step, the processing speed remains efficient. Using three memory pathways results in only an $18\%$ slowdown and an $8\%$ increase in GFlops, while delivering a 4.5-point improvement in performance.
@@
1 76.3 73.0 79.5 1$\times$ 844.1
2 80.1 76.7 83.5 0.93$\times$ 878.2
3 80.8 77.1 84.5 0.82$\times$ 912.3
4 80.7 77.0 84.5 0.75$\times$ 946.4
@@
4.4 Additional Results on VOT Benchmarks
To further evaluate the effectiveness of SAM2Long, we also apply it to three widely used single-object tracking benchmarks, namely LaSOT (Fan et al., 2019b), LaSOText (Fan et al., 2021), and GOT-10k (Huang et al., 2019). Without whistles and bells, we first generate an initial mask from the bounding box provided for the first frame using the SAM2 Mask decoder. Subsequently, all subsequent masks are transformed into bounding box format by identifying the minimum and maximum coordinates of their non-zero indices. For comparison, we include recent state-of-the-art methods specifically designed for VOT, including SAMURAI (Yang et al., 2024) and DAM2SAM (Videnovic et al., 2024).
Table 9: Performance comparison of SAM2Long with state-of-the-art VOT methods on LaSoT, LaSoText, and GOT-10k benchmarks. The results are reported in terms of AUC (Area Under the success rate Curve) for LaSoT and LaSoText, and AO (Average Overlap) for GOT-10k.
[…]
KeepTrack (Mayer et al., 2021) 67.1 48.2 -
TOMP (Mayer et al., 2022) 68.5 - -
DropTrack (Wu et al., 2023a) 71.8 52.7 75.9
SeqTrack (Chen et al., 2023) 72.5 50.7 74.8
MixFormer (Cui et al., 2022) 70.1 - 71.2
GRM-256 (Gao et al., 2023b) 69.9 - 73.4
ROMTrack (Cai et al., 2023) 71.4 51.3 74.2
OSTrack (Ye et al., 2022) 71.1 50.5 73.7
DiffusionTrack (Luo et al., 2024) 72.3 - 74.7
ODTrack (Zheng et al., 2024) 74.0 53.9 78.2
LORAT (Lin et al., 2025) 75.1 56.6 78.2
SAM2.1 (Ravi et al., 2024) 70.0 56.9 80.7
DAM2SAM (Videnovic et al., 2024) 75.1 60.9 81.1
SAMURAI (Yang et al., 2024) 74.2 61.0 81.7
SAM2Long (ours) 73.9 60.9 81.1
Based on the strong foundation of SAM2.1, we achieve competitive performance with state-of-the-art VOT methods. Specifically, our method, SAM2Long, delivers results that are comparable to the performance of SAMURAI and DAM2SAM, two recent methods specifically optimized for VOT.
@@ 5 Conclusion
In this paper, we introduce SAM2Long, a training-free enhancement to SAM 2 that alleviates its limitations in long-term video object segmentation. By employing a constrained tree memory structure with object-aware memory modulation, SAM2Long effectively mitigates error accumulation and improves robustness against occlusions, resulting in a more reliable segmentation process over extended periods. Extensive evaluations on six VOS benchmarks demonstrate that SAM2Long consistently outperforms SAM 2, especially in complex video scenarios. Notably, SAM2Long achieves up to a 5-point improvement in $\mathcal{J}\&\mathcal{F}$ scores on challenging long-term video benchmarks such SA-V and LVOS without requiring additional training or external parameters.